OpenStudy (joftheworld):

A missile is launched from the ground. Its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After one second, the missile is 130 feet in the air; after 2 seconds it is 240 feet in the air. Find the height, in feet, of the missile after 11 seconds in the air.

OpenStudy (cwrw238):

h(x) = ax^2 + bx + c as its fired from the ground we can assume c = 0 so we have 130 = a(1)^2 + 1b 240 = a(2)^2 + 2b solve this system of equations for a and b to get h(x) in terms of x

OpenStudy (cwrw238):

then plug in x = 11

OpenStudy (anonymous):

I really can't figure out this exact same question. Is there any way you could expand a little?